OpenStudy (anonymous):

How do you find the values of y if : x + (y-1)i = -1 + 3i , x = -1 ? x - 2y + (2x - y)i = 3i if x = 0 (1-i)x +(1+i)y = i (2-3i)x - (1+4i)y = 3 + i ?

OpenStudy (anonymous):

oh and if I didnt write x= ? for some then I have to find x too I think if x's value needs to be found

OpenStudy (anonymous):

You want to set one side to 0 then simplify as much as you can. first simplify then subistitute -1 in for x and solve. x+(y-1)i = -1+3i, x = -1 x+yi-1i = -1 + 3i -1 + yi - 1i =-1=3i //add 1 to both sides -1+1+yi-1i = 3i //subtract 3i from both sides -3i-1+1+yi-1i=0 //now simplify -4i + yi =0 // now subtract yi from both sides -4i = -yi //divide i from both sides -4i/i = -y //cross out i -4 = -y // convert signs 4 = y

OpenStudy (anonymous):

How do I solve x - 2y + (2x - y)i = 3i ?